OpenStudy (anonymous):

If Rob flips 3 coins what is the chance that they will come up all tails? Answer: a. 1/2 b. 1/4 c. 1/6 d. 1/8 e. 1/16

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }=?\] got it?

OpenStudy (hba):

@surdawi I warned you -_-

OpenStudy (hba):

I told you not to answer questions directly

OpenStudy (anonymous):

d(1/2)^3